# How Students Use AI Feedback Matters: Experimental Evidence on Physics Achievement and Autonomy

### v1

@@ 4.1.4 Results on autonomy (self-regulated learning)
According to Table 2, using self-regulated learning level as the dependent variable, interventions A and B produced marginally significant negative effects of 0.279 (p = 0.091) and 0.379 (p = 0.071) standard deviations, respectively, for the full sample. Analysis of sub-samples of the top, medium, and bottom thirds students revealed that GAI-powered hint recommendation and workbook answer and explanation recommendation significantly reduced the post-test self-regulated learning levels of high-scoring students by 0.477 (p < 0.05) and 0.973 (p < 0.05) standard deviations, respectively, whereas no significant effects were observed in the other sub-samples. Furthermore, no significant difference in effect sizes was detected between the two interventions.
@@ 4.2 Experiment 2: autonomous usage of GAI-powered on-demand help
To examine baseline equivalence among the three experimental groups, one-way analyses of variance (ANOVA) were conducted on pretest academic performance and pretest self-regulated learning scores for the full sample as well as for each achievement subgroup. For the full sample, results indicated no significant differences among the three groups in pretest academic performance, F(2, 263) = 0.56, p = .571, or in pretest self-regulated learning, F(2, 263) = 1.21, p = .301.
Similarly, within each academic achievement subgroup (high-, medium-, and low-achieving), no significant differences were observed among the three groups in either pretest academic performance (high-achieving: F(2, 84) = 2.56, p = .083; medium-achieving: F(2, 86) = 1.29, p = .280; low-achieving: F(2, 87) = 0.39, p = .677) or pretest self-regulated learning (high-achieving: F(2, 84) = 1.08, p = .344; medium-achieving: F(2, 86) = 1.16, p = .317; low-achieving: F(2, 87) = 1.94, p = .149).
@@ 4.2.4 Results on autonomy (learner autonomy)
Further analysis, which separated autonomy questionnaire responses into two theoretical dimensions—technical-psychological orientation (see Table 5) and political-social beliefs (see Table 6)—revealed that the negative impact of the Group D intervention was exclusively concentrated within the technical-psychological dimension (whole sample d = -0.299, p < 0.05; lower-achieving subgroup d = -0.549, p < 0.05), which aligns closely with self-regulated learning behaviors. In contrast, the structural relational belief dimension was unaffected. This result indicates that in this experiment the threat of using GAI on-demand help services on autonomy stayed in the frame of behaviors and cognition without interfering relational beliefs.
@@ 5.1 The value of GAI-powered hint versus traditional workbook explanations
In contrast, higher-achieving students engaged deeply with the feedback, including a thorough digestion of hints and additional practice problems provided by the system. In some cases, they even developed new ways of using the intervention tools—such as creating personalized notebooks from their feedback sheets. These findings suggest that when students are given the freedom to request resources and decide how to use them, those with stronger prior abilities are better equipped to take advantage of such opportunities. This observation aligns with established theories on self-regulated learning and echoes findings from interventions targeting various learner populations, which have shown that autonomy-supportive environments tend to benefit students who already possess higher levels of self-regulation (Deci & Ryan, 2000; Perry et al., 2006).
@@ 5.4 Implications, limitation and future directions
Taken together, these findings point toward a design principle for AI-powered educational interventions: provide more compulsory support for lower-achieving students, while offer greater autonomy and on-demand assistance to higher-achieving students. This differentiated approach may maximize benefits while minimizing risks to student autonomy.
The question of whether more advanced technology necessarily leads to better educational outcomes warrants scrutiny. When the present study’s findings on diminished learner autonomy in certain subgroups are considered alongside Stadler et al. (2024)’s evidence of reduced deep thinking and recent research on decreased agency (Zhang, 2024), a clear conclusion emerges: without thoughtful implementation, the integration of intelligent technologies into educational settings may be detrimental rather than beneficial to the cultivation of learners’ higher-order capacities. From Mencius (Chan, 2008) to Dewey (1938), the enduring wisdom that “adversity and challenge are essential to meaningful growth” serves as a timely warning. This calls for a nuanced approach to educational innovation, one that prioritizes the long-term development of learner autonomy over the uncritical adoption of technological advancement.
@@
Second, the study did not consistently employ the same questionnaire throughout. Initially, a self-regulated learning (SRL) questionnaire was used, later replaced with an autonomy questionnaire as the researchers’ understanding of the relevant constructs and measurement tools evolved. While this developmental process did not adversely affect the present study, employing a consistent set of instruments would have enhanced the reliability of the empirical results.

### v2

@@ 4.1.4 Results on autonomy (self-regulated learning)
According to Table 2, using SRL level as the dependent variable, interventions A and B produced marginally significant negative effects of 0.279 (p = 0.091) and 0.379 (p = 0.071) standard deviations, respectively, for the full sample. Analysis of sub-samples of the top, medium, and bottom thirds students revealed that GAI-powered hint recommendation and workbook answer and explanation recommendation significantly reduced the post-test SRL levels of high-scoring students by 0.477 (p < 0.05) and 0.973 (p < 0.05) standard deviations, respectively, whereas no significant effects were observed in the other sub-samples. Furthermore, no significant difference in effect sizes was detected between the two interventions.
@@ 4.2 Experiment 2: autonomous usage of GAI-powered on-demand help
To examine baseline equivalence among the three experimental groups, one-way analyses of variance (ANOVA) were conducted on pretest academic performance and pretest SRL scores for the full sample as well as for each achievement subgroup. For the full sample, results indicated no significant differences among the three groups in pretest academic performance, F(2, 263) = 0.56, p = .571, or in pretest SRL, F(2, 263) = 1.21, p = .301.
Similarly, within each academic achievement subgroup (high-, medium-, and low-achieving), no significant differences were observed among the three groups in either pretest academic performance (high-achieving: F(2, 84) = 2.56, p = .083; medium-achieving: F(2, 86) = 1.29, p = .280; low-achieving: F(2, 87) = 0.39, p = .677) or pretest SRL (high-achieving: F(2, 84) = 1.08, p = .344; medium-achieving: F(2, 86) = 1.16, p = .317; low-achieving: F(2, 87) = 1.94, p = .149).
@@ 4.2.4 Results on autonomy (learner autonomy)
Further analysis, which separated autonomy questionnaire responses into two theoretical dimensions—technical-psychological orientation (see Table 5) and political-social beliefs (see Table 6)—revealed that the negative impact of the Group D intervention was exclusively concentrated within the technical-psychological dimension (whole sample d = -0.299, p < 0.05; lower-achieving subgroup d = -0.549, p < 0.05), which aligns closely with SRL behaviors. In contrast, the structural relational belief dimension was unaffected. This result indicates that in this experiment the threat of using GAI on-demand help services on autonomy stayed in the frame of behaviors and cognition without interfering relational beliefs.
@@ 5.1 The value of GAI-powered hint versus traditional workbook explanations
In contrast, higher-achieving students engaged deeply with the feedback, including a thorough digestion of hints and additional practice problems provided by the system. In some cases, they even developed new ways of using the intervention tools—such as creating personalized notebooks from their feedback sheets. These findings suggest that when students are given the freedom to request resources and decide how to use them, those with stronger prior abilities are better equipped to take advantage of such opportunities. This observation aligns with established theories on SRL and echoes findings from interventions targeting various learner populations, which have shown that autonomy-supportive environments tend to benefit students who already possess higher levels of self-regulation (Deci & Ryan, 2000; Perry et al., 2006).
@@ 5.4 Implications, limitation and future directions
Findings of this study point toward a design principle for AI-powered educational interventions: provide more compulsory support for lower-achieving students, while offer greater autonomy and on-demand assistance to higher-achieving students. This differentiated approach may maximize benefits while minimizing risks to student autonomy.
The results on usage patterns, achievement and autonomy reveal a complex yet coherent picture: the educational effectiveness of GAI systems hinges not only on their content but, more critically, on how they are used and the extent to which the pattern of use aligns with student characteristics. In terms of achievement improvement, GAI-powered problem-solving hints show beneficial effects for low-achieving students under compulsory usage of personalized recommendations, whereas they support high-achieving students more effectively under autonomous, on-demand help conditions. However, regarding the development of autonomy, these same patterns reverse: the former harms high-achievers’ autonomy, and the latter negatively affects that of low-achievers.
These findings align with and expand upon previous research, such as Bastani et al. (2024)’s finding that student benefits more from interacting with agents that provide scaffolded prompts rather than direct answers, and Yan et al. (2025)’s demonstration that proactive GAI agents, which initiate questions, produce more sustained positive achievement gains than agents awaiting student input. Such evidence resonates with concerns raised by Stadler et al. (2024) and Darvishi et al. (2024), who caution that GAI-powered cognitive scaffolds diminish deep thinking and agency. In contrast, other studies—such as Du et al. (2024), who reported gains in self-regulation from recommending SRL strategies, and Li et al. (2025), who found metacognitive prompts effective in promoting SRL—highlight the potential for GAI to enhance autonomy when properly designed. Additionally, the present study examines one of the most widely implemented applications of GAI in education: problem-solving hints, demonstrating that GAI-powered cognitive support may enhance performance in exams without GAI assistance while undermining autonomy, depending on the usage pattern and learner profile. In light of the comparisons with others’ findings, the present study puts forward a critical insight: in the context of GAI for education, pedagogical patterns and models of usage matter.
@@
Second, the study did not consistently employ the same questionnaire throughout. Initially, a SRL questionnaire was used, later replaced with an autonomy questionnaire as the researchers’ understanding of the relevant constructs and measurement tools evolved. While this developmental process did not adversely affect the present study, employing a consistent set of instruments would have enhanced the reliability of the empirical results.